# SF-MMCN:A Low Power Re-configurable Server Flow Convolution Neural Network Accelerator

### v1

@@ Abstract
Convolution Neural Network (CNN) accelerators have been developed rapidly in recent studies. There are lots of CNN accelerators equipped with a variety of function and algorithm which results in low power and high speed performances. However, the scale of a PE array in traditional CNN accelerators is too big, which costs the most energy consumption while conducting multiply and accumulation (MAC) computations. The other issue is that due to the advance of CNN models, there are enormous models consist of parallel structures such as residual block in Residual Network (ResNet). The appearance of parallel structure in CNN models gives a challenge to the design of CNN accelerators owing to impacts on both operation and area efficiency. This study proposed SF-MMCN structure. The scale of PE array in proposed designs is reduced by pipeline technique in a PE. Proposed SF structure successfully make proposed SF-MMCN operate in high efficiency when facing parallel structures in CNN models. Proposed design is implemented with TSMC 90nm technology on VGG-16 and ResNet-18 environments. The performance of proposed design achieves 76% energy saving, 55% area saving and increases operation and are efficiency 9.25 times and 4.92 times respectively.
@@
Keywords Convolution Neural Network (CNN) $\cdot$ accelerators $\cdot$ parallel structure $\cdot$ area efficient $\cdot$ low power circuit
1 Introduction
@@
With the advanced of deep learning, the applications of Convolution Neural Network (CNN) have sprung up in recent years. The concept of CNN accelerators has emerged in a variety of architectures. However, due to the large number of parameters and the complexity of images skyrocket with the development of deep learning, the requirement of low power, high speed in CNN accelerator has been significant gradually. In order to reduce the scale of features in convolution computation, some studies Ramasamy et al. (2019); Mahdiani et al. (2009); Dalloo et al. (2018); Raghuram and Shashank (2022), proposed approximate convolution circuits, which directly decrease the computation resources by capturing partial data from input feature and weight data, the complexity of convolution would decrease distinctly. Moreover, because of the small size of the input data, the critical path of multipliers and the full adders become shorter, which would save the area and the power of the computation. And the timing cost by CNN can accelerate apparently. Since the contents of conventional multiplier are full adders, there have been many assembles of structures of multipliers nowadays, which have been same as approximate multiplier Joshi et al. (2023); Hashemi et al. (2015); Hammad et al. (2021); Liu et al. (2020a). In other words, implementing the approximate multiplier by improving the basic logic unit of the conventional multiplier becomes the most intuitive method. There are three main features of traditional multipliers:1) partial-product generate, 2) partial-product accumulate and 3) final addition. Generally, one of the serious problems of multipliers is that it wastes some times to accumulate the partial-product and final accumulate because the limited of the hardware. Consequently, compressors  Tonfat and Reis (2012); Sayadi et al. (2023); Strollo et al. (2020) reduce the partial-product and the critical path without any influence of circuit performances.
@@
Although approximate computation unit reduce the complexity of convolution apparently, the accuracy lost has become another serious issue of CNN accelerators. Due to the application of image recognition, accuracy lost is one of the most fatal factor of CNN accelerators Chen et al. (2016); Ansari et al. (2017); Yang et al. (2018); Ahmadi et al. (2021); Xu et al. (2020a); Liu et al. (2020b); Lu et al. (2022); Mo et al. (2021); Zeng et al. (2021); Xu et al. (2020b). Therefore, some researches try to focus on re-configurable structure of CNN accelerators instead of modifying algorithm of MAC computation. The CNN accelerator proposed in Ahmadi et al. (2021) is able to complete the convolution in different shape of kernel. And the dataflow of Ahmadi et al. (2021) pre-computes the sub-output and stores into memories in order to complete data reusing in $3\times 3$ convolution. When conducting the $1\times 1$ convolution, the architecture in Ahmadi et al. (2021) switched the convolution dimension from width to channel at a time. There are 195 PEs in Ahmadi et al. (2021) divided in 65 columns in top structure, which increase the throughput of the circuit. However, since the allocation of the PE array, the efficiency will decrease when the size of input features or weight data are not divided by 3 or 65.
@@
Due to bit-width in software is 64 bits, most of studies adapt fix-point precision technique to conduct hardware computation before the occurrence of quantization. However, by the lost data from fix-point data and data truncation between layers result in accuracy loss. Quantization trys to reduce the bit-width dramatically, but the accuracy is likely to maintain less than traditional method. Therefore, some researches adapt the quantization function into CNN accelerators. The implementation of this technique is exemplified in Mo et al. (2021), which introduces the Quantized Network-Acceleration Processor (QNAP) structure. In this innovative design, weight data, upon undergoing quantization, can be seamlessly assembled with other quantized weight data. Consequently, the PE structure in Mo et al. (2021) departs from conventional designs. With the quantization of weight data, Mo et al. (2021) reduces power consumption apparently.
@@
There are two CNN accelerator architecture between traditional structure and quantized structure, Figure 1. CNN accelerators with quantization technique require a compensate or quantization unit of hardware implementation, as indicated in Figure 1 (a). Due to lower bit-width data after quantization, the smaller area required of a PE array. On the other hand, traditional structure cost larger hardware area owing to elimination of a compensate or quantization unit Figure 1 (b).
@@
Another challenge in CNN accelerators is that due to the rapidly advanced of CNN models, there are lots of CNN models consist of parallel structures such as residual and bottleneck block. Ahmadi et al. (2021), Zeng et al. (2021) and Mo et al. (2021) implement the CNN accelerators with ability of operating parallel CNN models. However, to address parallel CNN structure, there are two methods of CNN accelerators in recent studies, parallel and series, both strategies result in different drawbacks of hardware implementations. For the parallel method, it would utilize more hardware to complete MAC operation due to large number of input data compared to the series method. However, the series strategy takes more time due to dividing residual blocks into several convolution layers. Therefore, it’s significant to strike the balance or get rid of both short-coming of two strategies mentioned above.
@@
This paper proposed a enhance version of Hsu et al. (2023), named Server Flow Multi-Mode Convolution Neural Network Unit (SF-MMCN). Proposed SF-MMCN achieves high power efficiency and area efficiency under parallel CNN models. With internal pipeline and SF structure, proposed SF-MMCN is able to feast in a variety of environments and overcome drawbacks result from the series and parallel strategies discussed before. Proposed SF-MMCN contributes to low power structure, small hardware area and high operation efficiency.
2 Definitions and Background
@@
As mentioned above, this paper proposed a CNN accelerator based on Multi-Mode Convolution Neural Network Unit (MMCN) in Hsu et al. (2023), Figure 2. MMCN proposed a computation core which supports convolution, max-pooling and dense function. All these functions share the same hardware so as to increase the utilization of the circuit. However, there are two serious issues of MMCN:
@@
  • Low efficiency on parallel CNN structure. The MMCN adopted series strategy when facing parallel CNN structure. Owing to operate each layer in series, MMCN takes much more time to complete the whole CNN model. However, the more executing time result in the more power consumption, which is against of the target of low power CNN accelerator.
@@
  • Large memory usage caused by the large scale of input data in parallel CNN structure. There isn’t any data reuse structure or technique in MMCN. As mentioned in Mo et al. (2021), data transmission between core and memories spend the most power of a chip. Therefore, although MMCN performs low computation power consumption, total power if hardware still not decreases apparently due to the huge requirement of memories.
Figure 2: MMCN structure in Hsu et al. (2023).
@@
Mo et al. (2021) proposed a dataflow inspired by Alwani et al. (2016) nad Mo et al. (2019). This strategy re-arranges parallel structure to series, similar to series method introduced before. However, it adds the pipeline technique between each layer. In single operation cycle, Mo et al. (2021) is able to execute each convolution in pipeline, Figure 3. Therefore, after conducting a parallel structure, take a residual block as an example, final outputs can be generated to output buffers. With this technique, Mo et al. (2021) speeds up the operation latency effectively. However, due to modified series strategy, the data transmission of data and memories still maintain high. The proposed SF structure in this paper reduce the power of data transmission apparently, which will be introduced in detail in next chapter.
Figure 3: Residual pipeline mode in Mo et al. (2021).
@@
3 Proposed Server Flow Multi-Mode Convolution Neural Network Unit
3.1 PE design
@@
A PE in SF-MMCN is showed in Figure 4. The most apparent modification of a PE is the selection between normal convolution output and convolution output with residual block. If the SF-MMCN is under residual mode, the MAC outputs will enter a adder to finish convolution computing with residual outputs. The MAC outputs will bypass to output registers to complete the normal convolution. A PE in SF-MMCN also equipped pipeline technique by a counter, which can enhance the throughput of whole structures.
@@
Proposed PE is also equipped with zero gate unit which is able to detect input image data. If input image data is zero, zero gate unit will turn off a multiplier and skip MAC operation to avoid redundant energy consumption. With zero gate unit, proposed PE can reduce additional power dramatically.
Figure 4: A PE in proposed-2 structure.
3.2 Pipeline Technique
@@
As depicted in Figure 4, a pipeline technique is integrated into each individual PE within the proposed SF-MMCN structure. Diverging from conventional CNN accelerators, including those in PE arrays like Zeng et al. (2021), the SF-MMCN structure optimizes chip area by incorporating the pipeline technique into each PE. Unlike the traditional approach of passing MAC temporary outputs sequentially from one PE to the next, the MMCN structure efficiently bypasses these computation paths through the utilization of the pipeline technique. To enhance the overall utilization and configuration of the proposed SF-MMCN structure, this study strategically employs counters to facilitate the implementation of the pipeline function.
3.3 Server Flow
Figure 5: Server Flow (SF) structure.
@@ 3.3 Server Flow
As previously discussed, within the PE array of the SF-MMCN architecture, one PE is assigned the responsibility of transmitting data to the other eight PEs, as illustrated in Figure 5. The dataflow during this process is demonstrated when a PE is passing data. During the computation of a series CNN model, the PE is typically set into idle mode to minimize power consumption. However, when confronted with a parallel CNN structure, SF initiates the corresponding computations, as depicted in Figure 6 (a), Figure 7 (a) and Figure 7 (b).
@@
During normal convolution operations in SF-MMCN, PE_9 is deactivated. The datapath for this scenario, along with the corresponding CNN structure, is illustrated in Figure 6. The MAC output, denoted as MAC1_9, swiftly enters a multiplexer, depicted by the red line in Figure 6 (a), to seamlessly complete the normal convolution process. While PE_9 may be considered redundant in this particular function, its deactivation contributes to the overall efficiency of the PE array, allowing for high-efficiency multi-mode operation through selective mode choices.
Figure 6: The dataflow of SF while executing.
During the execution of the residual function in SF-MMCN, without the presence of a residual convolution block, PE_9 becomes active. The hardware and software datapath for this scenario is depicted in Figure 7 (a) and (b). In this operation, designed for residual function without a residual convolution block, Figure 7 (a), the objective is to accumulate previous convolution outputs and the MAC outputs from PE_1 to PE_8. Consequently, PE_9 is specifically tasked with transmitting the previous convolution output to a multiplexer controlled by the testbench (Mode select 1). The output from PE_9 then enters an adder situated near the rest of the PEs, where it accumulates with a MAC output (MAC1_9) to complete the computation. This configuration streamlines the process, facilitating the efficient execution of residual functions within the SF-MMCN architecture.
Figure 7: (a) The closer look of SF structure under normal residual structure. (b) The closer look of SF structure under residual structure with a convolution layer.
Figure 8: The dataflow of proposed SF on/off.
When SF-MMCN is engaged in the residual function with a residual convolution block, the output path of PE_9 transitions to its dedicated Multiply-Accumulate (MAC) output, as illustrated in Figure 7 (b). This output, along with the MAC outputs from PE_1 to PE_8, is directed to an adder to accomplish convolution. In this capacity, PE_9 functions akin to a server, tasked with preparing the residual convolution output. Importantly, considering the shape of weights in residual convolution, SF can efficiently set up the preparation of residual convolution output in a timely manner, ensuring synchronization with the completion of MAC computations by the other eight PEs.
@@ 3.4 Operation Flow
Figure 9: Illustration of the input feature and corresponded executed PEs in proposed SF-MMCN.
@@
Figure 8 shows the waveform of SF-MMCN, PO represents partial output. Single convolution can be complete in 10 cycles if the width and height of a filter is $3*3$. Once input and weight data enter SF-MMCNs, the MAC operation executes immediately. After 9 input features and weights finish loading, proposed SF-MMCN will take 1 cycle to get the final convolution outputs. When encountering residual structure, proposed SF-MMCN consumes the same cycles as normal convolution. As mentioned above, PE_1 to PE_8 compute all MAC data of normal convolution, and PE_9 would deliver or compute the previous convolution output and MAC operation respectively during the same cycles. Therefore, proposed SF-MMCN is able to complete the residual structure without redundant computation cycles.
The proposed SF-MMCN performs high throughput due to SF structure. Owing to four SF-MMCNs in implementation architecture Figure 13, the shape of output data by PE_1 to PE_8 is $3\times 3\times 8$, which means $width\times height\times channel$ and the output shape of PE_9 is $1\times 1\times 8$, Figure 9. The value of channel equals to the number of the SF-MMCN in the implementation. The batch size of proposed SF-MMCN is 1 because of high speed convolution requirement of CNN accelerators Ahmadi et al. (2021). In other words, due to batch size is 1, proposed SF-MMCN is able to provide decision result in real time.
Figure 10: The shape of input features of proposed SF-MMCN under different conditions.
As introduced above, Figure 10 shows the input data shape of proposed under different environment. When encountering series structure, the shape of output is the same as input shape with zero-padding, $3*3*8$. While facing parallel CNN structures, the throughput increases because of PE_9 execute the input feature of parallel block. The allocation of a input feature map is indicated in Figure 10 with different colors.
Another problem is that proposed can address the input features with small size. Take $2\times 2$ input map as an instance, the PE array will separate into two parts to complete two channels of input data, Figure 11. Therefore, proposed SF-MMCN reduces the potential of redundant circuit under different complexities of environment.
@@
Figure 11: The dataflow of proposed SF-MMCN on a small size input feature.
3.5 Data Reuse

### v2

# SF-MMCN: Low-Power Sever Flow Multi-Mode Diffusion Model Accelerator

@@ Abstract
Generative Artificial Intelligence (AI) has become incredibly popular in recent years, and the significance of traditional accelerators in dealing with large-scale parameters is urgent. With the diffusion model’s parallel structure, the hardware design challenge has skyrocketed because of the multiple layers operating simultaneously. Convolution Neural Network (CNN) accelerators have been designed and developed rapidly, especially for high-speed inference. Often, CNN models with parallel structures are deployed. In these CNN accelerators, many Processing Elements (PE) are required to perform parallel computations, mainly the multiply and accumulation (MAC) operation, resulting in high power consumption and a large silicon area. In this work, a Server Flow Multi-Mode CNN Unit (SF-MMCN) is proposed to reduce the number of PE while improving the operation efficiency of the CNN accelerator. The pipelining technique is introduced into Server Flow to process parallel computations. The proposed SF-MMCN is implemented with TSMC 90-nm CMOS technology. It is evaluated with VGG-16, ResNet-18, and U-net. The evaluation results show that the proposed SF-MMCN can reduce the power consumption by 92%, and the silicon area by 70%, while improving the efficiency of operation by nearly 81 times. A new FoM, area efficiency (GOPs/$mm^{2}$) is also introduced to evaluate the performance of the accelerator in terms of the ratio throughput (GOPs) and silicon area ($mm^{2}$). In this FoM, SF-MMCN improves area efficiency by 18 times (18.42).
Index Terms:
@@
accelerators, area efficiency, Convolution Neural Network (CNN), diffusion model, parallel structure, low-power circuit.
I Introduction
@@
With the advanced of deep learning, the applications of Convolution Neural Network (CNN) have sprung up in recent years. The concept of CNN accelerators has emerged in a variety of architectures. However, due to the large number of parameters and the complexity of images skyrocketing with the development of deep learning, the requirement of low-power, high speed in CNN accelerator has been significant gradually. To reduce the scale of features in convolution computation, some studies [1, 2, 3, 4], the proposed approximate convolution circuits, which directly decrease computation resources by capturing partial data from input feature and weight data, the complexity of convolution would decrease distinctly. Moreover, because of the small size of the input data, the critical path of multipliers and the full adders becomes shorter, which would save the area and power of the computation. And the timing cost by CNN can accelerate. Since the contents of a conventional multiplier are full adders, there have been many assembles of structures of multipliers nowadays, which have been the same as the approximate multiplier [5, 6, 7, 8]. In other words, implementing the approximate multiplier by improving the basic logic unit of the conventional multiplier becomes the most intuitive method. There are three main features of traditional multipliers: 1) partial-product generation, 2) partial-product accumulation, and 3) final addition. Generally, one of the serious problems of multipliers is that it takes some time to accumulate the partial product and the final accumulate because it is limited by the hardware. Consequently, compressors  [9, 10, 11] reduce the partial product and the critical path without any influence on circuit performances.
@@
Although approximate computation units reduce the complexity of convolution, the accuracy lost has become another serious issue of CNN accelerators. Due to the application of image recognition, accuracy loss is one of the most fatal factors of CNN accelerators [12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Therefore, researchers try to focus on the reconfigurable structure of CNN accelerators instead of modifying the algorithm of MAC computation. The CNN accelerator that was proposed in [15] can complete the convolution in different shapes of the kernel. The data flow of [15] precomputes the sub-output and stores it in memories to complete the reuse of the data in $3\times 3$ convolution. When performing the $1\times 1$ convolution, the architecture in [15] changed the dimension of the convolution from width to channel at a time. There are 195 PEs in [15] divided into 65 columns in the top structure, increasing the throughput of the circuit. However, since the PE array is allocated, the efficiency will decrease when the size of the input features or the weight data are not divided by 3 or 65.
@@
Due to the bit-width in software being 64 bits, most studies adopt the fix-point precision technique to conduct hardware computation before the occurrence of quantization. However, the lost data from fix-point data and data truncation between layers result in accuracy loss. Quantization tries to reduce the bit-width dramatically, but the accuracy is likely to be less than the traditional method. Therefore, researchers adapt the quantization function into CNN accelerators. The implementation of this technique is exemplified in [19], which introduces the Quantized Network-Acceleration Processor (QNAP) structure. In this innovative design, weight data, upon undergoing quantization, can be seamlessly assembled with other quantified weight data. Consequently, the PE structure in [19] departs from conventional designs. With the quantization of weight data, [19] reduces power consumption.
@@
There are two CNN accelerator architectures between traditional structure and quantified structure, Figure 1. CNN accelerators with quantization technique require a compensation or quantization unit of hardware implementation, as indicated in Figure 1 (a). Due to lower bit-width data after quantization, a smaller area is required of a PE array. On the other hand, the traditional structure cost a larger hardware area owing to the elimination of a compensate or quantization unit Figure 1 (b).
@@
Another challenge in CNN accelerators is that due to the rapid advancement of CNN models, there are lots of CNN models consisting of parallel structures such as residual and bottleneck blocks. [15], [19], and [20] implement the CNN accelerators with the ability to operate parallel CNN models. However, to address parallel CNN structure, there are two methods of CNN accelerators in recent studies, parallel, and series, both strategies result in different drawbacks of hardware implementations. The parallel method would utilize more hardware to complete MAC operation due to the large number of input data compared to the series method. However, the series strategy takes more time due to dividing residual blocks into several convolution layers. Therefore, it’s significant to strike a balance or get rid of both shortcomings of the two strategies mentioned above.
Furthermore, given the trend of the diffusion model [22] for generative AI, U-net [23], as one of the most popular models for the de-noise diffusion model, consists of many parallel structures. Since generative AIs are responsible for dealing with enormous data. What’s more, the applications of generative AI mostly require real-time characteristics such as translation and use as a personal secretary. Consequently, accelerating the AI model with a parallel structure has become the most necessary challenge in recent years.
@@
This paper proposed an enhanced version of [24, 25], named Sever Flow Multi-Mode Convolution Neural Network Unit (SF-MMCN). The proposed SF-MMCN achieves high power efficiency and area efficiency under parallel CNN models. With an internal pipeline and SF structure, the proposed SF-MMCN can feast in a variety of environments and overcome drawbacks resulting from the series and parallel strategies discussed before. By adjusting the mode of hardware, the proposed structure supports not only the residual block but also the time parameter process in U-net. The proposed SF-MMCN contributes to a low power structure, small hardware area, and high operation efficiency.
II Definitions and Background
@@
As mentioned above, this paper proposed a CNN accelerator based on a Multi-Mode Convolution Neural Network Unit (MMCN) in [24], Figure 2. MMCN proposed a computation core that supports convolution, max-pooling, and dense functions. All these functions share the same hardware to increase the utilization of the circuit. However, there are two serious issues of MMCN:
@@
  • Low efficiency on parallel CNN structure. The MMCN adopted a series strategy when facing a parallel CNN structure. Owing to operating each layer in series, MMCN takes much more time to complete the whole CNN model. However, more execution time results in more power consumption, which is against the target of a low-power CNN accelerator.
@@
  • Large memory usage is caused by the large scale of input data in parallel CNN structure. There isn’t any data reuse structure or technique in MMCN. As mentioned in [19], data transmission between core and memories has the most power of a chip. Therefore, although MMCN performs low computation power consumption, the total power of hardware still does not decrease apparently due to the huge requirement for memories.
Figure 2: MMCN structure in [24].
@@
In [19], a dataflow inspired by [26] and [27] was proposed. This strategy re-arranges parallel structure to series, similar to the series method introduced before. However, it adds the pipeline technique between each layer. In a single operation cycle, [19] can execute each convolution in the pipeline. Therefore, after conducting a parallel structure, take a residual block as an example, final outputs can be generated to output buffers. With this technique, [19] speeds up the operation latency effectively. However, due to the modified series strategy, the data transmission of data and memories remained high. The proposed SF structure in this paper reduces the power of data transmission apparently, which will be introduced in detail in the next section.
For the diffusion model, the training flow is completed by adding noise to the original for a great scale of iteration while prediction is the opposite of training. The noise has to be removed from the previous feature map input to achieve the final clear output. This action is called ”de-noise” in the diffusion model. The accelerator has to conduct thousands or even millions of times to get the output figure. To achieve this goal, power consumption and latency are the most difficult issues for traditional accelerators. Therefore, this paper proposed a structure that is equipped with the ability to conduct de-noise computation under fast, low-power figures.
Figure 3: De-noise dataflow of diffusion model [22].
@@
III Sever Flow Multi-Mode Convolution Neural Network Unit
The design and implementation of SF-MMCN will be discussed in this section.
III-A PE design
@@
A PE in SF-MMCN is shown in Figure 4. The most apparent modification of a PE is the selection between normal convolution output and convolution output with residual block. If the SF-MMCN is under the residual mode, the MAC outputs will enter an adder to finish convolution computing with residual outputs. The MAC outputs will bypass to output registers to complete the normal convolution. A PE in SF-MMCN also equipped pipeline technique by a counter, which can enhance the throughput of whole structures.
@@
The proposed PE is also equipped with a zero gate unit which can detect input image data. If input image data is zero, the zero gate unit will turn off a multiplier and skip MAC operation to avoid redundant energy consumption. With a zero gate unit, the proposed PE can reduce additional power dramatically.
Figure 4: A PE in the proposed-2 structure.
III-B Pipeline Technique
@@
As depicted in Figure 4, a pipeline technique is integrated into each PE within the proposed SF-MMCN structure. Diverging from conventional CNN accelerators, including those in PE arrays like [20], the SF-MMCN structure optimizes chip area by incorporating the pipeline technique into each PE. Unlike the traditional approach of passing MAC temporary outputs sequentially from one PE to the next, the MMCN structure efficiently bypasses these computation paths through the utilization of the pipeline technique. To enhance the overall utilization and configuration of the proposed SF-MMCN structure, this study strategically employs counters to facilitate the implementation of the pipeline function.
Figure 5: Sever Flow (SF) structure.
Figure 6: (a) The dataflow of SF while executing. (b) The closer look of SF structure under normal residual structure. (c) The closer look of SF structure under residual structure with a convolution layer.
III-C Sever Flow
@@
As previously discussed, within the PE array of the SF-MMCN architecture, one PE is assigned the responsibility of transmitting data to the other eight PEs, as illustrated in Figure 5. The data flow during this process is demonstrated when a PE is passing data. During the computation of a series CNN model, the PE is typically set into idle mode to minimize power consumption. However, when confronted with a parallel CNN structure, SF initiates the corresponding computations, as depicted in Figure 6 (a) and Figure 6 (b).
@@
During normal convolution operations in SF-MMCN, PE_9 is deactivated. The datapath for this scenario, along with the corresponding CNN structure, is illustrated in Figure 6 (a). The MAC output, denoted as MAC1_9, swiftly enters a multiplexer, depicted by the red line in Figure 6 (a), to seamlessly complete the normal convolution process. While PE_9 may be considered redundant in this particular function, its deactivation contributes to the overall efficiency of the PE array, allowing for high-efficiency multi-mode operation through selective mode choices.
During the execution of the residual function in SF-MMCN, without the presence of a residual convolution block, PE_9 becomes active. The hardware and software datapath for this scenario is depicted in Figure 6 (b) an d (c). In this operation, designed for residual function without a residual convolution block, Figure 6 (b), the objective is to accumulate previous convolution outputs and the MAC outputs from PE_1 to PE_8. Consequently, PE_9 is specifically tasked with transmitting the previous convolution output to a multiplexer controlled by the testbench (Mode select 1). The output from PE_9 then enters an adder situated near the rest of the PEs, where it accumulates with a MAC output (MAC1_9) to complete the computation. This configuration streamlines the process, facilitating the efficient execution of residual functions within the SF-MMCN architecture.
Figure 7: The dataflow of the proposed SF on/off.
When SF-MMCN is engaged in the residual function with a residual convolution block, the output path of PE_9 transitions to its dedicated Multiply-Accumulate (MAC) output, as illustrated in Figure 6 (c). This output, along with the MAC outputs from PE_1 to PE_8, is directed to an adder to accomplish convolution. In this capacity, PE_9 functions akin to a server, tasked with preparing the residual convolution output. Importantly, considering the shape of weights in residual convolution, SF can efficiently set up the preparation of residual convolution output promptly, ensuring synchronization with the completion of MAC computations by the other eight PEs.
@@
Figure 8: The illustration of SF-MMCN with different size of input feature map.
III-D Operation Flow
@@
Figure 7 shows the waveform of SF-MMCN, PO represents partial output. Single convolution can be completed in 10 cycles if the width and height of a filter are $3*3$. Once input and weight data enter SF-MMCNs, the MAC operation executes immediately. After 9 input features and weights finish loading, the proposed SF-MMCN will take 1 cycle to get the final convolution outputs. When encountering residual structure, the proposed SF-MMCN consumes the same cycles as normal convolution. As mentioned above, PE_1 to PE_8 compute all MAC data of normal convolution, and PE_9 would deliver or compute the previous convolution output and MAC operation respectively during the same cycles. Therefore, the proposed SF-MMCN can complete the residual structure without redundant computation cycles.
@@
Figure 9: Illustration of the input feature and corresponding executed PEs in the proposed SF-MMCN.
The proposed SF-MMCN performs high throughput due to the SF structure. Owing to four SF-MMCNs in implementation architecture Figure 18, the shape of output data by PE_1 to PE_8 is $3\times 3\times 8$, which means $width\times height\times channel$ and the output shape of PE_9 is $1\times 1\times 8$, Figure 9. The value of the channel equals the number of the SF-MMCN in the implementation. The batch size of the proposed SF-MMCN is 1 because of the high-speed convolution requirement of CNN accelerators [15]. In other words, due to the batch size being 1, the proposed SF-MMCN can provide decision results in real-time.
Figure 10: The shape of input features of the proposed SF-MMCN under different conditions.
As introduced above, Figure 10 shows the input data shape of the proposed under different environments. When encountering a series structure, the shape of the output is the same as the input shape with zero padding, $3*3*8$. While facing parallel CNN structures, the throughput increases because PE_9 executes the input feature of the parallel block. The allocation of an input feature map is indicated in Figure 10 with different colors.
Another problem is that the proposed can address the input features with a small size. Take $2\times 2$ input map as an instance, the PE array will separate into two parts to complete two channels of input data, Figure 11. Therefore, the proposed SF-MMCN reduces the potential of redundant circuits under different complexities of the environment.
Figure 11: The dataflow of the proposed SF-MMCN on a small size input feature.
Consequently, Figure 8 demonstrates the datapath of the proposed SF-MMCN with a normal and small input feature map under both common CNN structures and residual blocks. The green and yellow squares are responsible for series and parallel structure respectively.
As the address of an input feature map of the residual block is not continued, the address can be generated by shifters. The corresponding input data pass into the PE_9 directly to finish MAC computing. According to Figure 8, it’s straightforward forward in implement eight input data of a residual block that can be well-distributed to the rest of PEs to complete MAC operation. Therefore, SF-MMCN doesn’t waste any additional cycle waiting for MAC output from PE_1 to PE_8.
As described in Figure 11, the control unit separates the PE array into two parts to complete computations. When facing residual structure, the PE_9 of the proposed SF-MMCN can complete two channels of computations before the MAC outputs are delivered from other PEs due to the same pixel number of both series block and parallel block.
Figure 12: The dataflow of PE_9 on a normal/small size input feature.
Figure 12 indicated the dataflow of PE_9 under the different sizes of input. Before the whole system receives the final answer, PE_9 passes the corresponding MAC data to each PE under normal input image size conditions. On the other hand, when the condition changes to a small input image size, channel N and channel N+1 in Figure 12 is the distribution separated by the control unit. At the beginning of four cycles, PE_9 is responsible for the computations of channel N, the rest of the four cycles are the computations of the MAC data of channel N+1. Therefore, due to the hardware and timing distribution of the proposed SF-MMCN, there aren’t any occurrences of redundant circuits and cycles.
III-E Diffusion model
Figure 13: The U-net structure.
Figure 13 shows the U-net which is implemented in this paper. There are lots of blocks in the model. Each block consists of two convolution layers and one dense layer which is responsible for the computation of time parameters. Moreover, the hardware of each block is the same, which means that the proposed SF-MMCN can support multiple computations in the same operation cycle. The dataflow of a block of U-net is indicated in Figure 14. There are 4 groups of a single block, which distributes the U-net block into a dense layer (Block 1), a convolution layer with an activation function (Block 2), a convolution layer without an activation function (Block 3), and final logic computation (Block 4).
Figure 14: The block distribution of U-net.
According to the block distribution, the dataflow with timing is described by Figure 15 and Figure 16. When the computation starts (T0 in Figure 15), Figure 16 PE_1 to PE_8 (orange) are responsible for the convolution with activation function (ReLu). Meanwhile, PE_9 conducts a time parameter dense layer. From T1 to T2 in Figure 15, PE_1   PE_8 operates the other convolution layer and combines the final result through final logic computation.
Figure 15: The operation timing distribution of the proposed SF-MMCN.
Figure 16: The allocation of the proposed structure among different blocks.
By introducing the implementations of the proposed SF-MMCN, the key to the proposed SF is PE_9. Since the rest of the PEs are recommended to fulfill the main connection of models, PE_9 is assisted to deal with the parallel structure such as residual and time parameter operator. This architecture seems not quite complex but can complete difficult, deep learning models. Therefore, the proposed structure is equipped with great configurable features.
III-F Data Reuse